Question 4:

Current and projected free cash flows for ABC Inc. are shown here. Growth is expected to be constant after 2021, and the weighted average cost of capital is 11%. What is the horizon (continuing) value at 2022 if growth from 2021 remains constant?

Actual

Projected

2019

2020

2021

2022

Free cash flow (millions of dollars)

$606.820

$667.500

$707.547

$750.000
Answer #1
Solution:
Horizon (continuing) value at 2022 15,900
Working Notes:
Notes: Horizon value is the present value of all the FCF after that period ,at end of calculated period
First of all we calculate constant growth rate
constant growth rate (g) = (FCF 2022 - FCF 2021)/FCF 2021
constant growth rate (g) = (750 - 707.547)/707.547
constant growth rate (g) =0.06
constant growth rate (g) =6%
Horizon (continuing ) Value at 2022 = FCF 2022 ( 1 + g)/(WACC - g)
Horizon (continuing ) Value at 2022 = 750 ( 1 + 6%)/(11% - 6%)
Horizon (continuing ) Value at 2022 = 795/5%
Horizon (continuing ) Value at 2022 = 15,900
Please feel free to ask if anything about above solution in comment section of the question.